OpenStudy (anonymous):

I need help solving this equation using the elimination method: 8s+12t=16, 6s+9t=12

OpenStudy (compassionate):

8s + 12t = 16 6s + 9t = 12 You want to add the bottom and top together.

OpenStudy (anonymous):

I'm am math illiterate. So, I would add and make it 14s+21t=28?

OpenStudy (campbell_st):

why... they are the same equation... divide the 1st equation by 4 divide the 2nd equation by 3 what do you have...?

OpenStudy (anonymous):

2s+3t=4

OpenStudy (compassionate):

You want to s or t terms to cancel out in either equation. I can multiply the entire 2nd equation by -4 and the entire 1st equation by 3 in order to cancel out the s-terms. 3(8s + 12t = 16) -4(6s + 9t = 12) 24s + 36t = 48 -24s - 36t = - 48 Now add both the equations to one-another.

OpenStudy (campbell_st):

yep... so the equations have an infinite number of solutions....

OpenStudy (compassionate):

Campbell, the equations have no solutions...

OpenStudy (campbell_st):

as you can see with @Compassionate method.... adding the equations will give 0 = 0

OpenStudy (campbell_st):

they are the same line... if you graphed them both... they sit on top of each other... so have an infinite number of points in common... but that's just my interpretation.

OpenStudy (campbell_st):

no solution would mean the lines are parallel... never intersect..

OpenStudy (compassionate):

You're correct, Campbell!

OpenStudy (campbell_st):

phew... that was lucky

OpenStudy (anonymous):

I don't know either way. I'm in an online class and the teacher is from a foreign country and didn't explain it very well! Thank you!!!

OpenStudy (compassionate):

Ha ha.

OpenStudy (compassionate):

Jewels, if you need anymore help just message an Ambassador or member. We're here for ya'. ;)

OpenStudy (anonymous):

I do have another one. I really need tutoring on all of this. This one may be done the same way but he didn't specify if he wanted us to use elimination or substitution. 4x-6y=3, 6x-4y=-3

OpenStudy (compassionate):

It would be best to use elimination 4x - 6y = 3 6x - 4y = -3 Okay, so first you need to say, "Can I multiply my x or y terms to get them to cancel out on another?" Well, 4x and 6x have a common factor. Do you know what it is?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2(4x-6y=3) -2(6x-4y=-3) ??

OpenStudy (compassionate):

No no, I mean, at what point do 4 and 6 reach the same number? For example, 2 and 5 have the same multiple of 10, because they're both divisble by each other.

OpenStudy (anonymous):

24 so it would be: 6(4x-6y=3) -4(6x-4y=-3) Right?

OpenStudy (compassionate):

Right! So we have 24x - 6y = 3 -24x - 6y = - 3 Nwo we add the terms and our xs wil cancel and we can solve for y.

OpenStudy (anonymous):

6y=-9 6y/6=-9/6 y=-1.5 right?

OpenStudy (compassionate):

Correct. Now take -1.5 and plug it into any of the two equations to solve for x.

OpenStudy (anonymous):

4x-6(-1.5)=3 4x+9=3 4x+(9-9)=(3-9) 4x=-6 is that right so far?

OpenStudy (compassionate):

Yup. Now divide by 4.

OpenStudy (anonymous):

x=-1.5 y=-1.5

OpenStudy (compassionate):

Right! so those are your x and y intercepts. Now if you want to find the slope just use the slope formula.

OpenStudy (anonymous):

That one I do know. lol. Thank you so much! I think I'm beginning to understand it a little be easier. :)

OpenStudy (compassionate):

Yup! And that's called the Elimination Method! :)

OpenStudy (anonymous):

Thank you!